solve for the resistance of a 100m long tungsten wire which has a circular cross sectional area with a diameter of 0.1mm solutions
1
Expert's answer
2020-02-27T10:10:18-0500

The resistance is


"R=\rho\frac{l}{A}=\rho\frac{4l}{\pi d^2}"

"R=5.60\cdot 10^{-8}\frac{4(100)}{\pi (1\cdot 10^{-4})^2}=713\ \Omega"